How many moles of specified particles are present in a sample of each of the following substances if each sample weighs 7.00 g?
a. N2O molecules
b. NO2 molecules
c. P4O10 molecules
d. V atoms
1
Expert's answer
2020-10-08T13:58:47-0400

(a) N2O molecules

RAM of N=14, O=16

RMM = (14 x2) + 16 = 44g/mol


Moles "=\dfrac{mass}{RMM} = \dfrac{7g}{44g/mol}"


= 0.159moles


(b) NO2 molecules


RMM = 14 + (16 x 2) = 46g/mol


Moles "= \dfrac{7g}{46g/mol} = 0.152 moles"


(c) P4O10 molecules


RAM of P=31, O=16


RMM = (31 x 4) + (16 x 10) = 284g/mol


Moles "= \dfrac{7g}{284g/mol} = 0.0246 moles"


(d) V atoms

RAM of V = 50.9g/mol


Moles "= \dfrac{7g}{50.9g} = 0.139 moles"